OpenStudy (anonymous):

Tom is 4 years older than Jerry. Nine years ago, Tom was 5 times as old as Jerry was then. How old is each now?

OpenStudy (callisto):

Let x be the age of Tom now. ''Tom is 4 years older than Jerry.'', what is the age of Jerry now? @jungbunny

OpenStudy (anonymous):

wait but how do you set up the equation?

OpenStudy (callisto):

Just do it step by step. Can you tell me if Tom is x years old now, how old is Jerry?

OpenStudy (anonymous):

hi, can you take jerry's present age be x then tom is four years older that implies tom's age is (x+4) after that talk about nine years before age of both of them.. are you following..?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

nine yearss before, jerry should be of age nine less than his present age , which implies jerry is (x-9) years old then..

OpenStudy (anonymous):

let's talk about tom, his age must also be 9 less than his present age.. all followed ..?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

now, they have in question given a relation relating tom's and jerry's age nine years before... saying that "Tom was 5 times as old as Jerry was then"

OpenStudy (anonymous):

so you simply write it in equation form !!

OpenStudy (anonymous):

okay, tom's age nine years before = (x+4)-9 do you agree..?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

here's the equation ! tom's age nine year before = five times( jerry's age nine years before) this means (x+4)-9 = x_9 solve this...

OpenStudy (anonymous):

sorry that's minus and not underscore in above equation !

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

hopefully you understood !!

OpenStudy (anonymous):

you have to use distributive proberty right?

OpenStudy (anonymous):

for what ? you know how to solve lenear equations , right ?

OpenStudy (anonymous):

not really...i don't remeber how to.

OpenStudy (anonymous):

i'm sorry the equatin is (x+4)-9= 5(x-9) i missed a nine ! and yes darling your right use distributive property !

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

:) you're welcome